OpenStudy (anonymous):

(-4x^2 -5x-1)(4x^2 -6x-2) show work

OpenStudy (anonymous):

-4x^2 (4x^2 -6x -2) - 5x(4x^2 -6x -2) -1(4x^2 -6x -2) Distribute and add

OpenStudy (anonymous):

@Compassionate ?

OpenStudy (kainui):

9= 3*3= (1+2)(1+1+1)= 1(1+1+1)+2(1+1+1)= 1(1+1+1)+1(1+1+1)+1(1+1+1)= 3+3+3= 9 Just sort of playing around with it sort of makes it intuitively obvious.

OpenStudy (anonymous):

you have to use foil @kainui